# Parameter-free prediction of irradiation defect structures in tungsten at room temperature using stochastic cluster dynamics Sicong He1, Brandon Schwendeman2, George Tynan2, Jaime Marian1,3

### v1

@@ 4.2 Discussion on literature on heavy ion irradiation of W
The majority of heavy-ion irradiation studies in W have been performed at temperatures near the operational window for fusion applications, 500∘C. Under such conditions, monovacancies are mobile, leading to the formation of vacancy clusters and voids of various sizes depending on irradiation conditions. Irradiations with 18-MeV self-ions at 500 and 800∘C up top 0.2 dpa by Hwang et al. [80] revealed both dislocation loops and voids, with average sizes of, respectively, 3.1 and 1.01.5 nm. These vacancy clusters are at or below the visibility threshold in TEM, which hampers our ability to extract statistically meaningful counts of these defects. This limitation can be addressed using positron annihilation spectroscopy (PAS) (e.g., refs. [81, 32]), which can detect down to monovacancies. Early PAS tests have determined with precision the temperature for Stage III recovery, which was found to be between 523 and 573 K [81]. More recent PAS results of irradiations with 18-MeV self-ions to 0.0085 dpa indicate that approximately 98
In terms of room temperature tests, Wang et al. [82] irradiated W foils with 6.0-MeV Cu ions up to a damage level of 0.6 dpa, finding a microstructure dominated by dislocation loops with an average size of 8.0 nm () at concentrations greater than m-3. These are one to two orders of magnitude larger than what we report here, consistent with a total irradiation dose 75 times higher than ours. These authors also pointed out that the loop microstructure saturates after 0.2 dpa of irradiation, which is not seen here to our lower doses, as demonstrated by the increasing trends in Fig. 6(b). In contrast, their finding that the main vacancy defects were monovacancies and small clusters is consistent with our calculations. A separate work involving 20-MeV self-ion irradiation of polycrystal W with an average grain size of 40 microns [77] reports the accumulation of visible defect clusters in the amounts shown in Fig. LABEL:sia-concentration. In the range of doses common to our study and that in ref. [77], and despite the differences in the material microstructure and the irradiation conditions, there is very good quantitative agreement between our simulations (for TEM-visible defects) and the experimental data. In particular, our simulations provide an almnost exact match of the time evolution of the cluster population observed experimentally.
Finally, being a mean-field approach, our study does not discriminate between different Burgers vectors for the dislocation loops. Recent experiments using of 800-keV Kr ions at 400∘C suggest a 50/50 partition between and Burgers vectors [49]. Likewise, TEM specimens irradiated with 1.2-MeV W ions up to 0.017 dpa at room temperature indicate a similar loop coexistence [83]. While these results may be at odds with some atomistic simulations, which predict a clear general preponderance of loops [54, 84, 85], better agreement is found if only vacancy-type loops are considered.
4.3 Comparison between model predictions and experimental data
The simulated defect cluster distributions show excellent overall agreement with the experimental measurements across nearly four orders of magnitude in concentration. The XRDS measurements provide sufficient resolution for clusters with . As such, there are no experimental data against which to compare the model predictions directly in the range. For SIA defects (Fig. LABEL:sia-scatter) in the as-irradiated condition, the populations of clusters with show an increasing discrepancy between the measured data as the cluster size decreases. This discrepancy is mitigated in the annealed condition, suggesting that the combined effect of thermal coarsening and monomer dissociation during a prolonged annealing period is important at low cluster sizes. Additionally, part of the deviation can be attributed to the use of closed-form size-dependent cross-sections for dislocation loops in the interpretation of XRDS signals [31], which are known to misrepresent the structure of SIA clusters with [86].
For the vacancy cluster subpopulation, Fig. LABEL:vac-scatter, the disparity between the calculations and measurements is more pronounced at small sizes than for SIA defects, even in the annealed condition. In addition to the signal resolution factor discussed for SIA clusters, the XRDS cross sections used for small vacancy clusters are also obtained assuming ‘loop’-like defect structures. However, as discussed in the previous subsection, PAS experiments have conclusively established the globular nature of vacancy clusters with . Indeed, recent measurements by Hu et al. [32] obtained from experiments conducted at 500∘C with 20-MeV ions up to 0.02 dpa deomonstrate high concentrations of clusters with .
What we can say with certainty is that, for both SIA and vacancy clusters with , the agreement between the model predictions and the experimental data is outstanding, which adds a high level of confidence to the underlying physical principles behind the simulation models. Future work will include scenarios corresponding to different irradiation conditions, particularly temperature, particle energy, and particle mass (to control the PKA energy spectrum), with the ultimate goal of projecting to fusion energy conditions in leading conceptual designs.
@@ 5 Conclusions
We have assembled a number of fundamental physics features that have recently emerged in the literature into a numerical model for irradiation damage accumulation in tungsten. The model includes athermal SIA diffusion along rectilinear paths, atomistically-generated Frenkel-pair insertion correlations and cluster size distributions, and large sampling volumes to ensure that an adequate sampling of the high-energy tail of the PKA energy distribution is effectuated.
The model relies heavily on pre-computed atomistic properties and contains no adjustable parameters.
The model has been validated mainly through a series of XRDS experiments of self-ion-irradiated single-crystal W at room temperature. The experiments were designed as to facilitate comparison with the model by simplifying the irradiation conditions and eliminating microstructural complexities.
The model is also in very good agreement with other independent works on W irradiated at room temperature.
Model predictions are in excellent agreement with XRDS measurements of vacancy and SIA cluster size distributions at the end of a 0.085-dpa irradiation stage followed by a long annealing at room temperature. This agreement holds over a large range of cluster sizes for both self-interstitial and vacancy defect types.
Finally, the model provides quantitative predictions in the missing cluster size range outside the resolution limit of XRDS, thus complementing experimental data beyond what experiments can access.
@@ Acknowledgments
The authors acknowledge financial support from the U.S. Department of Energy, Office of Fusion Energy Sciences (DOE-FES) under contracts DE-SC0018410 (S.H. and J.M.) and DE-SC0022528 (B.S. and G.T.). This work used computational and storage services associated with the Hoffman2 Cluster which is operated by the UCLA Office of Advanced Research Computing’s Research Technology Group. The authors acknowledge the use of facilities and instrumentation at the UC Irvine Materials Research Institute (IMRI) which is supported in part by the National Science Foundation through the UC Irvine Materials Research Science and Engineering Center (grant No. DMR- 2011967).
@@ Appendix A Brief review of the stochastic cluster dynamics method
The accumulation of defects during irradiation is simulated using the stochastic cluster dynamics (SCD) model. SCD is a stochastic variant of the mean-field rate theory technique which relies on stochastic sampling of the underlying master equation for defect cluster evolution [87, 88]. Instead of deterministically solving exceedingly large sets of partial differential equations (PDE) of the concentrations of defects, as in standard MFRT, SCD evolves an integer-valued defect population in a finite material volume , thus avoiding exponential growth in the number of ODEs. This makes SCD ideal to treat problems where the dimensionality of the cluster size space is high, e.g., when multispecies simulations are of interest [89, 22, 90].
@@
where the set represents the reaction rates for defect insertion, thermal dissociation, and annihilation at sinks, and binary reactions occurring inside . Subindices and refer to all distinct defect species present in the system, whereas the subindex refers to the type of sinks available to absorb defects (‘d’ for dislocations, ‘gb’ for grain boundaries, ‘ppt’ for precipitates). The subindices in the expression satisfy and (reflecting dissociation by monomer emission and association of complementary clusters, respectively). Equation (5) can be straightforwardly extended to partial differential equations to capture diffusion due to defect concentration gradients [91].
Equation (5) is solved using the residence-time algorithm by sampling, selecting, and executing events from the set of rates [87]. The volume is arbitrary, although its minimum value is subjected to the numerical stability criterion:
@@
where and are the diffusivity and the lifetime of mobile species , . For consistency with the experimental conditions provided above, here we consider vacancies, self-interstitial atoms, and any cluster combination thereof.
As a practical note, under certain conditions the volume can be rescaled for improved computational efficiency, leading to an accelerated defect cluster evolution. A quantitative demonstration of volume rescaling applied to a selected group of defect clusters is provided in ref. [92]. For more information about the model, the reader is referred to past publications from the authors [87, 88, 91].
@@ A.1 Defect insertion rates
The defect insertion rates, in eq. (5), are obtained for each defect type by multiplying the volumetric insertion rates ( in Fig. 1) times the simulation volume . In turn, the rates are set by the ion insertion rate, defined in eq. (4), followed by sampling recursively until the total ion energy expended on lattice damage, i.e., 2.65 MeV, is reached (see Sec. 2.2 for more details)444Note that using the ASTM-recommended value for the threshold displacement energy of W of =90 eV, this amount of energy would result in the creation of Frenkel pairs according to the NRT formalism [93] when an efficiency of is used. However, this is known to grossly overpredict the number of defects actually created during irradiation [94].. During each sampling, defect distributions obeying eqs. (2) and using the scaling laws extracted from Fig. 4(a) are generated, and an array of species is created. These species are then added to the simulation volume and let to evolve according to eqs. (5).
In order to appropriately sample the entire range of cluster sizes, set by the histograms in Fig. 4(a), it is desirable to produce PKA with energies greater than 50 keV. Without them, the large clusters observed experimentally with would not be created in the simulations. However, due to the low probability of producing PKA with those energies (it is worth keeping in mind that the average PKA energy in the present irradiations is 1 keV), large numbers of PKA samplings must be instantiated. This can only be achieved by maintaining sufficiently high values of (eq. (4)), which is directly proportional to the simulation volume . Thus, a crucial part of our model is ensuring that is kept above a minimum limit that guarantees adequate sampling of and the cluster size power laws.
@@ A.2 Defect energetics
We also consider thermal dissociation of SIA and vacancy clusters by emission of monomers. As indicated in Fig. 1, the dissociation rate scales with the factor , where is the binding energy of a monomer (monovacancy or single SIA) to a cluster of size . Atomistic calculations of SIA and vacancy clusters up to exist in the literature [18, 14, 84]. Specifically, here we use the DFT data by Becquart et al. [95], as listed in Table 3. For larger sizes, the binding energies are obtained as:
where is the formation energy of a cluster of size . Here we assume that vacancy clusters adopt a globular shape for , whose formation energies are well captured using the capillary approximation, i.e., [14, 91]. The expression for the binding energies in this size range is given in Table 3. Finally, we assume that vacancy clusters with and SIA clusters with appear as prismatic loops with energies given by the expression [84, 85]:
from which, after operating with eq. (9), we obtain the final expression for the binding energy of loops:
and are numerical coefficients given for each type of loop in Table 3.

### v3

@@ 4.3 Discussion on literature on heavy ion irradiation of W
A large body of heavy-ion irradiation studies in W have been performed at temperatures near the operational window for fusion applications, 500∘C. Under such conditions, monovacancies are mobile, leading to the formation of vacancy clusters and voids of various sizes depending on the irradiation conditions. Irradiations with 18 MeV self-ions at 500 and 800∘C up top 0.2 dpa by Hwang et al. HWANG2016430 revealed both dislocation loops and voids, with average sizes of, respectively, 3.1 and 1.01.5 nm. These vacancy clusters are at or below the visibility threshold in TEM, which hampers our ability to extract statistically meaningful counts of these defects. This limitation can be addressed using positron annihilation spectroscopy (PAS) (e.g., refs. debelle2008first, hu2025new), which can detect down to monovacancies. Consistent with the aforementioned TEM results, recent PAS measurements by Zavašnik et al. zavavsnik2025microstructural of W crystals irradiated with 10.8 MeV self-ions to 0.02 and 0.2 dpa at 800 K similarly report the presence of larger vacancy clusters comprised of 2550 point defects (approximately 1 nm in diameter) at elevated irradiation temperatures. Interestingly, despite the presence of larger clusters, new detailed PAS analyses of W irradiated with 18-MeV self-ions to 0.0085 dpa nevertheless indicate that approximately 98% of the total concentration of vacancy clusters consists of defects with hu2025new. Although these irradiations were performed at, or around, 500∘C, the presence of large nm dislocation loops and large densities of small vacancy clusters are consistent with our results in Fig. 6(a).
Recent studies of ion-irradiated W employing a combination of TEM and PAS Wang_2024 have demonstrated that the larger TEM-visible voids observed in the studies referenced above at elevated temperatures result from the thermal evolution of the defect population due to mobilized monovacancies. PAS studies of ion-irradiated W near room temperature, below the Stage III recovery temperature debelle2008first, Keys_1970, Wang_2024, have thus reported the presence of large densities of monovacancies and very small vacancy clusters, with no voids visible in TEM imaging Wang_2023, Wang_2024. Hollingsworth et al. Hollingsworth_2022 performed PAS measurements after 2-MeV self-ion irradiation to 0.0085 dpa at room temperature and reported a population of primarily monovacancies with an estimated concentration of m-3. This result agrees remarkably well with the modeling predictions and XRDS-based inference of the monovacancy density (see Fig. LABEL:vac-scatter in the previous subsection) for the higher-energy ions used here. With increasing irradiation dose, this same work noted a trend toward increasing contributions from small clusters to the measured spectra. Zavašnik et al. zavavsnik2025microstructural also performed PAS measurements of W single crystals irradiated at room temperature with 10.8 MeV self-ions, nearly identical to the conditions studied in this work. At both 0.02 and 0.2 dpa, they reported a population of primarily monovacancies and very small, vacancy clusters, in general agreement with the previous work and the modeling results presented here in Fig. 6(a).
TEM imaging for room temperature irradiations of W using 6.0 MeV Cu ions by Wang and co-workers wang2023defect, Wang_2023, Wang_2024 over a dose range of 0.05 to 0.6 dpa revealed a microstructure dominated by dislocation loops with an average size of 68 nm () at concentrations on the order of m-3. Despite the higher irradiation dose range in comparison to that studied here, this loop concentration is over an order of magnitude lower than the TEM-visible SIA cluster concentration in Fig. LABEL:sia-scatter (assumed to be dislocation loops and defined in the figure as ) at the end of the irradiation period, corresponding to 0.008 dpa. Nevertheless, Figure 6(c) illustrates that the absolute dislocation loop size distribution measured by Wang et al. Wang_2023 at the most comparable dose of 0.05 dpa is remarkably consistent with the results of this work for larger loop sizes nm. No loop nature distinction was made in their work, so the results are compared directly in Fig. LABEL:compa-size to the total cluster size distribution obtained by adding the SIA cluster and vacancy cluster concentrations in Figs. LABEL:sia-scatter and LABEL:vac-scatter for both the SCD modeling predictions and the XRDS measurements. The observed discrepancy for smaller loop sizes 6 nm is consistent with a detailed study comparing XRDS and TEM measurements of defect cluster size distributions in irradiated nickel by Olsen et al. Olsen_2016, which illustrates that the TEM resolution limit is not sharp. Instead, for clusters nm in diameter, the TEM measurements obtained using the bright field imaging method utilized by Wang et al. Wang_2023 were shown to increasingly undercount the defect cluster concentrations with decreasing size. Note that 6-nm loops in W correspond to , which is much larger than the smallest loop sizes that are often visible in TEM measurements that are undercounted due to resolution effects. Importantly, Wang et al. Wang_2023 also pointed out that the loop microstructure saturates after an irradiation dose of 0.2 dpa, which is not seen here due to our lower doses, as demonstrated by the monotonically increasing trends in Fig. 6(b). Finally, we note that their finding that the main vacancy defects were monovacancies and small clusters after irradiation at room temperature is consistent with the studies discussed above and the present calculations.
A separate body of work by Yi et al. YI2016105 involving 150-keV self-ion irradiations of W thin foils at room temperature reports the accumulation of TEM-visible dislocation loops in the amounts shown in Figure 6(c). These authors reported a predominance of vacancy loops at low doses ( 0.01 dpa) and room temperature, with a tendency towards equipartition as the dose increases ( 0.1 dpa). These observations are consistent with an accompanying study done using 2-MeV W-ion irradiations in semi-bulk tungsten specimens yi2015characterisation. However, establishing the defect nature of loops at low doses is based mostly on qualitative analysis of ‘black-and-white’ contrast of TEM micrographs, and on the erroneous assumption that SIA loops are formed by diffusion and agglomeration of individual SIA defects, and can therefore only appear at high doses after sufficient SIA production has taken place haussermann1972study. For that reason, here we make no assumptions about their nature and plot their concentrations directly in Figure LABEL:compa-size next to the total cluster size distributions for both the SCD modeling predictions and the XRDS measurements as described above. Despite the differing irradiation conditions, the figure shows remarkable agreement between the TEM-measured loop size distributions by Yi et al. YI2016105 and those obtained in this work for loops greater than 4 nm in diameter. The improved agreement at smaller loop sizes in comparison to the Wang et al. Wang_2023 TEM results is consistent with the higher resolution weak beam dark field imaging method used in the study by Yi et al. YI2016105, which Olsen et al. Olsen_2016 demonstrated undercounts defect clusters smaller than 34 nm in diameter.
Likewise, Fig. LABEL:compa-dose shows the buildup of visible loops (i.e., with diameters larger than 1.52.0 nm) with irradiation dose as predicted by the SCD simulations (SIA and vacancy loops combined) and those measured experimentally using in-situ TEM in the study by Yi et al. YI2016105. In the range of doses common to our study and theirs, there is approximately a factor of 20 difference in total concentrations. This is not surprising given the differences in the irradiation species, W specimens (i.e. thin foils versus bulk), and the unsharp nature of the TEM resolution limit discussed above. Nonetheless, our simulations provide an almost exact match of the time exponents (, 1.1) of the cluster population observed experimentally.
Finally, we discuss the nature and Burgers vectors of loops formed directly in displacement cascades in irradiated W. Prismatic loops may be of either self-interstitial or vacancy character, and can carry Burgers vectors of or . This gives rise to four distinct loop types that may form under irradiation and must therefore be distinguished. The distinction is important because the energetics and mobilities of each type can be drastically different. In particular, loops are highly glissile, while their counterparts are effectively immobile in most practical situations marian2003md. As well, as indicated in A.2, the dissociation energies of point defects from each loop type can also differ substantially gilbert2008structure, mason2019relaxation, which may have implications on their thermal stability during annealing. However, being a mean-field approach, SCD does not discriminate between different Burgers vectors for the dislocation loops. Recent experiments using 800-keV Kr ions at 400∘C suggest a 50/50 partition between and Burgers vectors ZHENG2020162. Similarly, TEM samples irradiated with 1.2-MeV W ions up to 0.017 dpa at room temperature indicate similar loop coexistence hasanzadeh2018three. Atomistic simulations generally predict a preponderance of loops byggmastar2025four, alexander2016initio, mason2019relaxation, although the observation of loops in MD simulations of high-energy cascades is not uncommon, and has been reported in some studies sand2013high, setyawan2015displacement, byggmastar2025four.
@@ 5 Conclusions
We have assembled a number of fundamental physics features that have recently emerged in the literature into a numerical model for the accumulation of irradiation damage in tungsten. The model includes atomistically-generated Frenkel-pair insertion correlations and cluster size distributions, athermal SIA diffusion along rectilinear paths, and large sampling volumes to ensure that an adequate sampling of the high-energy tail of the PKA energy distribution is effectuated.
The model contains no adjustable parameters, and its accuracy is thus derived from the incorporation of pre-computed atomistic properties and physics mechanisms primarily derived from detailed atomistic simulations.
The model was validated through a series of XRDS experiments of self-ion-irradiated single-crystal W at room temperature. The experiments were designed to facilitate comparison with the model by simplifying the irradiation conditions and eliminating microstructural complexities.
Model predictions are in excellent agreement with XRDS measurements of vacancy and SIA cluster size distributions at the end of a 0.008 dpa irradiation stage followed by a long annealing at room temperature. This agreement holds over a wide range of cluster sizes for both self-interstitial and vacancy defect types.
The model is also in excellent agreement with other independent works on W irradiated under similar conditions near room temperature, predicting large densities of monovacancies and small vacancy clusters consistent with numerous PAS studies and predicting absolute defect cluster distributions in the larger TEM-visible size range comparable to existing studies.
Finally, the model provides a comprehensive picture of the SIA and vacancy cluster size distributions in irradiated W, thus complementing the capabilities of disparate experimental techniques with differing sensitivities to specific cluster morphologies and sizes.
@@ Acknowledgments
The authors acknowledge financial support from the U.S. Department of Energy, Office of Fusion Energy Sciences (DOE-FES) under contracts DE-SC0018410 (S.H. and J.M.) and DE-SC0022528 (B.S. and G.T.). This work used computational and storage services associated with the Hoffman2 Cluster which is operated by the UCLA Office of Advanced Research Computing’s Research Technology Group. The authors acknowledge the use of facilities and instrumentation at the UC Irvine Materials Research Institute (IMRI), which is supported in part by the National Science Foundation through the UC Irvine Materials Research Science and Engineering Center (grant No. DMR- 2011967).
@@ Appendix A Brief review of the stochastic cluster dynamics method
The accumulation of defects during irradiation is simulated using the stochastic cluster dynamics (SCD) model. SCD is a stochastic variant of the mean-field rate theory (MFRT) technique, which relies on stochastic sampling of the underlying master equation for defect cluster evolution RN22, RN20. Instead of deterministically solving exceedingly large sets of partial differential equations (PDEs) of the concentrations of defects (as in standard MFRT), SCD evolves an integer-valued defect population in a finite material volume , thus avoiding exponential growth in the number of ODEs. This makes SCD ideal to treat problems where the dimensionality of the cluster size space is high, e.g., when multispecies simulations are of interest DUNN201643, MARIAN2012293, hu2025helium.
@@
where the set represents the reaction rates for, respectively, defect insertion; thermal dissociation and annihilation at sinks; and binary reactions occurring inside . Subindices and refer to all distinct defect species present in the system, whereas the subindex refers to the type of sinks available to absorb defects (‘d’ for dislocations, ‘gb’ for grain boundaries, ‘ppt’ for precipitates). The subindices in the expression satisfy and (reflecting dissociation by monomer emission and association of complementary clusters, respectively). Equation (5) can be straightforwardly extended to partial differential equations to capture diffusion due to defect concentration gradients Yu_2020.
Equation (5) is solved using the residence-time algorithm by sampling, selecting, and executing events from the set of rates RN22. The volume is arbitrary, although its minimum value is subjected to the numerical stability criterion:
@@
where and are the diffusivity and the lifetime of mobile species with . For consistency with the experimental conditions provided above, here we consider vacancies, self-interstitial atoms, and any cluster combination thereof. The number of clusters of type can be straightforwardly (and correctly) converted to a concentration by normalizing by the volume : .
As a practical note, under certain conditions the volume can be rescaled for improved computational efficiency, which leads to an accelerated defect cluster evolution. A quantitative demonstration of volume rescaling applied to a selected group of defect clusters is provided in ref. HE2024155325. For more information about the model, the reader is referred to past publications from the authors RN22, RN20, Yu_2020.
@@ A.1 Defect insertion rates
The defect insertion rates, in eq. (5), are obtained for each defect type by multiplying the volumetric insertion rates ( in Fig. 1) times the simulation volume . In turn, the rates are set by the ion insertion rate, defined in eq. (4), followed by sampling recursively until the total ion energy expended on lattice damage, i.e., 2.65 MeV, is reached (see Sec. 2.2 for more details)444Note that using the ASTM-recommended value for the threshold displacement energy of W of =90 eV, this amount of energy would result in the creation of Frenkel pairs according to the NRT formalism norgett1975proposed when an efficiency of is used. However, this is known to grossly overpredict the number of defects actually created during irradiation zinkle2023quantifying.. During each sampling, defect distributions obeying eqs. (2) and using the scaling laws extracted from Fig. 4(a) are generated, and an array of species is created. These species are then added to the simulation volume and allowed to evolve according to eqs. (5).
In order to appropriately sample the entire range of cluster sizes illustrated by the histograms in Fig. 4(a), it is desirable to produce PKA with energies greater than 100 keV, as, without them, the large clusters observed experimentally with would not be created in the simulations. However, due to the low probability of producing PKA with those energies (note that the average PKA energy in the present irradiations is 1 keV), large numbers of PKA samplings must be instantiated. This can only be achieved by maintaining sufficiently high values of (eq. (4)), which is directly proportional to the simulation volume . Thus, a crucial part of our model is ensuring that is kept above a minimum limit that guarantees adequate sampling of and the cluster size power laws.
@@ A.2 Defect energetics
We also consider thermal dissociation of SIA and vacancy clusters by emission of monomers. As indicated in Fig. 1, the dissociation rate scales with the factor , where is the binding energy of a monomer (monovacancy or single SIA) to a cluster of size . Atomistic calculations of SIA and vacancy clusters up to exist in the literature fikar2009molecular, becquart2010microstructural, alexander2016initio. Specifically, here we use the DFT data by Becquart et al. BECQUART201039, as listed in Table 3. For larger sizes, the binding energies are obtained as:
where is the formation energy of a cluster of size . Here we assume that vacancy clusters adopt a globular shape for , whose formation energies are well captured using the capillary approximation, i.e., becquart2010microstructural, Yu_2020. The expression for the binding energies in this size range is given in Table 3. Finally, we assume that vacancy clusters with and SIA clusters with appear as prismatic loops with energies given by the expression alexander2016initio, mason2019relaxation:
Combined with eq. (9), we obtain the final expression for the binding energy of loops:
where and are numerical coefficients given for each type of loop in Table 3.
| Species | [eV] |
| 2.12 | |
| 3.02 | |
| 3.60 | |
| 3.98 | |
| 4.27 | |
| 5.39 | |
| 0.04 | |
| 0.64 | |
| 0.72 | |
| 0.89 | |
| 0.72 | |